# Three-view Focal Length Recovery From Homographies

### v1

@@ 5.3 Robust Estimation of Focal Lengths
To estimate focal lengths from three images of planar scenes we utilize the LO-RANSAC framework [8] using a strategy inspired by [37]. We first extract triplet point correspondences from images (e.g. using [12, 34]). In each RANSAC iteration we sample 4 triplets from which we estimate $\mathbf{G}_{2}$ and $\mathbf{G}_{3}$ using DLT [20]. These matrices are used as inputs to solvers proposed in Section 5.1 and 5.2.
For each resulting real positive solution we use (7) to obtain $\mathbf{H}_{2}$, which is then decomposed into two possible poses $(\mathbf{R}_{2},\mathbf{t}_{2})$. We then use these poses and focal lengths to triangulate three of the sampled points in the first two views thus obtaining points $\mathbf{X}_{i1}$. We use the corresponding points in the third view to obtain $\mathbf{R}_{3}$ and $\mathbf{t}_{3}$ using the $\mathbf{P}3\mathbf{P}$ solver [14]. Note that it is possible to obtain $\mathbf{R}_{3}$ and $\mathbf{t}_{3}$ by decomposing $\mathbf{H}_{3}$ and using a single correspondence to obtain the scale of $\mathbf{t}_{3}$, but in practical experiments we found the approach using $\mathbf{P}3\mathbf{P}$ faster and more accurate. We score the generated models using pairwise Sampson error for each of the three image pairs. Whenever a new so-far-best model is found we perform local optimization using the Levenberg-Marquardt algorithm minimizing the pairwise Sampson error across the three pairs of views. In addition to the estimated focal lengths this strategy also produces the relative poses of all three cameras. We denote the proposed robust estimators as $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ for Case I, $\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ for Case II, $\mathbf{H}_{f\rho\rho}+\mathbf{P}3\mathbf{P}$ for Case III and $\mathbf{H}_{f\rho}+\mathbf{P}3\mathbf{P}$ for Case IV to distinguish them from the solvers. Our implementation is based on PoseLib [31].
@@ 6 Experiments
We perform extensive experiments on synthetic and real data to evaluate the performance of the proposed solvers and the robust focal length estimation strategy described in Section 5.3. We compare our solvers with several baselines using either pairwise or triplet correspondences. For pairwise correspondences, we consider the 6 point solver for relative pose with one unknown shared focal length [27] denoted as $f\mathbf{E}f$and its combination with the $4+1$ point plane and parallax solver [44] using the DEGENSAC framework [9], which we denote as $f\mathbf{E}f+\mathbf{P}\mathbf{P}$. We also evaluate the solver for one unknown focal length [5] denoted as $\mathbf{E}f$. When considering triplets, we use a strategy [37] of using a pairwise solver followed by triangulation of points and registration of the third camera using the $\mathbf{P}3\mathbf{P}$ solver [14]. (see Section 5.3 for more details). We denote the methods that work with triplet correspondences as $f\mathbf{E}f+\mathbf{P}3\mathbf{P}$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$, $\mathbf{E}f+\mathbf{P}3\mathbf{P}$ respectively.
Note that we do not compare with the solvers proposed in [23]. As visible from Table 1 these solvers are significantly slower than our solvers  [^2] The reported runtimes are runtimes of the original Matlab implementations of [23] and the Matlab + mex implementations of our solvers on I7-11700K CPU. and they return significantly more solutions that need to be tested inside RANSAC. As such they are not practical for real-world applications. Moreover, the solver for Case IV proposed in [23] returns solutions only to one of the two unknown focal lengths.
@@ 6.1 Synthetic experiments
$\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ (ours)$\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ (ours)$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$ [45]$f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$$\mathbf{E}f+\mathbf{P}3\mathbf{P}$$f\mathbf{E}f$ [27]$f\mathbf{E}f+\mathbf{P}\mathbf{P}$ [44]$\mathbf{E}f$ [5]
@@
Case I $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ ours 4 triplets 0.0439 0.1491 51.49 65.58 49.31
$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$ [45] 6 triplets 0.0503 0.2215 47.83 61.45 37.44
@@
$f\mathbf{E}f$ [27] 6 pairs 0.4793 0.7143 10.49 17.07 24.54
$f\mathbf{E}f+\mathbf{P}\mathbf{P}$ [44] 6 pairs 0.5166 0.7935 10.31 16.51 31.95
Case II $\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ ours 4 triplets 0.0611 0.2244 43.11 55.54 35.97
$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$ [45] 6 triplets 0.0692 0.2438 40.42 52.35 31.81
@@
$f\mathbf{E}f$ [27] 6 pairs 0.4623 0.7085 12.01 19.12 28.59
$f\mathbf{E}f+\mathbf{P}\mathbf{P}$ [44] 6 pairs 0.4968 0.7875 12.11 18.98 37.60
$\mathbf{E}f$ [5] 6 pairs 0.6827 0.6843 18.87 13.01 22.59
@@
$\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ (ours)$\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ (ours)$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$$f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$$\mathbf{E}f+\mathbf{P}3\mathbf{P}$
@@
The synthetic data is generated with $n=200$ 3D points visible by the three $1920\text{px}\times 1080\text{px}$ cameras with focal lengths sampled unfiormly $f_{g}\in\left[300,3000\right]$. The cameras are positioned randomly such that their views overlap. $n_{p}$ of the total $n$ points lie on a plane with random orientation. From the 3D points we create triplets with 75% inlier ratio by randomly shuffling 25% of the points. We add Gaussian noise with standard deviation of $\sigma$ to the pixel coordinates of points. We run all of the tested methods in PoseLib [31] for a fixed number of 100 iterations with the Sampson error threshold of 3 px. For methods that use only pairwise correspondences we consider only the first two views. For each configuration we generate 100 random scenes.
Fig. 3 shows the results of the synthetic experiments in which we vary both $\sigma$ and $n_{p}$. For Case II we also show how the error in the known focal length $\rho$ affects the accuracy of the methods (Fig. 3 (f)). When considering Case I, $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ shows significantly better accuracy than the remaining methods when a large portion of points lie on a single dominant plane (Fig. 3 (a)) even under high levels of noise (Fig. 3 (b)). This indicates a clear superiority of $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ when dealing with images of planar scenes. Similarly, for Case II, $\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ also shows improved focal length estimation accuracy when considering planar scenes, although the improvement over competing methods is not as significant. For fully planar scenes, as expected, the two-view solvers $f\mathbf{E}f$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ and $\mathbf{E}f$ fail to generate correct solutions (Fig. 3 (b),(e),(f)). For scenes with few of the plane points, e.g., scenes with $n_{p}/n\in\left[0.95,0.9\right]$, the $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ solver, i.e., the solver where the $4+1$ plane+parallax solver is utilized within DEGENSAC with $f\mathbf{E}f$ [44], outperforms the $f\mathbf{E}f$ solver (Fig. 3 (a)). This is not surprising since the $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ solver [44] was developed to handle scenes with dominant planes and some off-the plane points. The solvers that work with triplet correspondences, i.e. $f\mathbf{E}f+\mathbf{P}3\mathbf{P}$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$, $\mathbf{E}f+\mathbf{P}3\mathbf{P}$, outperform the two-view solvers.
@@ 6.2 Real-World Experiments
Since there are no suitable datasets for evaluating focal length recovery on planar scenes, we have collected a new dataset for evaluation which we will make publicly available. We have collected the dataset using 14 different cameras. For each camera we obtained ground truth focal lengths using the standard calibration method [46] with a checkerboard pattern. Using these cameras we have captured four indoor and two outdoor planar scenes from various positions resulting 1870 images in total. Sample images of the scenes are shown in Fig. 5. To obtain triplet correspondences we used SuperPoint features [12] with LightGlue matches [34]. More details are available in SM.
@@
Using the proposed solvers we may obtain multiple solutions. Therefore these solvers need to be used within a robust estimation framework such as RANSAC. Other simpler strategies such as voting [6] may fail due to the fact that more than one solution may be geometrically feasible.
Moreover, as mentioned above for Case III and IV focal lengths can not be distinguished using only planar points thus requiring some off-plane points during scoring. Note that this is a property of the problem and not the solvers. Problem with recovering one focal length for Case III was mentioned also in [23].
@@ 8 New Constraints
Here we provide details on the constraints introduced in Sec. 4 of the main paper. We show the steps for their derivation including Macaulay2 [17] code.
To derive the constraints relating the focal lengths and the elements of the matrix $\mathbf{Q}_{i}$, we first create an ideal $I$ [11] generated by the 12 polynomials extracted form the matrix equation Eq (16) in the main paper, i.e. the equation
$\begin{split}[\mathbf{n}]_{\times}\mathbf{Q}_{j}[\mathbf{n}]_{\times}^{\top}=s_{j}[\mathbf{n}]_{\times}[\mathbf{n}]_{\times}^{\top},\ j=2,3\\ \end{split}$ (26)
@@
In the next step, the unknown elements of the normal vector $\mathbf{n}$, i.e., $n_{x},n_{y},n_{z}$, and the scale factors $s_{2},s_{3}$ are eliminated from the generators of $I$ by computing the generators of the elimination ideal $J=I\cap\mathbb{C}[q_{21},\dots,q_{36}]$. Here, $q_{j.}$ are the entries of $\mathbf{Q}_{2},\mathbf{Q}_{3}$. These generators can be computed, for example, in the computer algebra software Macaulay2 [17] using the following code:
@@
$\mathbf{H}_{\rho ff}+\mathbf{P}3\mathbf{P}$ (ours)$\mathbf{H}_{\rho f}+\mathbf{P}3\mathbf{P}$ (ours)$f\mathbf{E}f+\mathbf{P}4\mathbf{P}f$$f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}4\mathbf{P}f$$\mathbf{E}f+\mathbf{P}4\mathbf{P}f$
@@ 10 Evaluation for Case III and Case IV
In this section we evaluate the robust estimators $\mathbf{H}_{\rho ff}+\mathbf{P}3\mathbf{P}$ for Case III and $\mathbf{H}_{f\rho}+\mathbf{P}3\mathbf{P}$ for Case IV which were presented in Sec.5.3 of the main paper. As baselines for comparison we use the combination of the $f\mathbf{E}f$solver [27] or its DEGANSAC [9] variant with the plane and parallax solver [44] together with the $\mathbf{P}4\mathbf{P}f$ solver [28] for Case III and the $\mathbf{E}f$ solver in combination with $\mathbf{P}4\mathbf{P}f$ solver [28] for Case IV.
@@ 10.1 Multiple Geometrically Valid Solutions
We observe that for both Case III and Case IV a single planar scene may result in multiple geometrically valid solutions. This makes it difficult to select a single correct solution for a given set of point correspondences. This leads to generally worse performance of the proposed solvers than for Case I and Case II. Note, that this is a feature of the problems and not the solvers. Problem with recovering one focal length for Case III was mentioned also in [23].
@@ 10.3 Real-World Experiments
In Tab. 3 we show the results comparing our method and the baseline approaches in three different variants. As the first variant we do not discard any solutions. For the second variant, we set range of acceptable FOVs by considering the prior for focal lengths used by the popular SfM software COLMAP [40] which is set as $f_{p}=1.2\text{max}(width,height)$, which corresponds to a field of view of $\sim 45^{\circ}$. To obtain a range we consider 30% increase or decrease in focal length, resulting in the acceptable field of view range from 35.5 to 61.5. As the last variant we use a range of 50 to 70. We chose this range since most commercially available phone cameras fall within it.
All variants were implemented in PoseLib [31]. We set the maximum epipolar threshold to 3 px. We used early termination with 0.9999 confidence and a minimum of 100 iterations and a maximum of 1000.

### v2

@@ 5.3 Robust Estimation of Focal Lengths
To estimate focal lengths from three images of planar scenes we utilize the LO-RANSAC framework [chum2003loransac] using a strategy inspired by [nister2004efficient]. We first extract triplet point correspondences from images (e.g. using [detone2018superpoint, lindenberger2023lightglue]). In each RANSAC iteration we sample 4 triplets from which we estimate $\mathbf{G}_{2}$ and $\mathbf{G}_{3}$ using DLT [hartley2003multiple]. These matrices are used as inputs to solvers proposed in Section 5.1 and 5.2.
For each resulting real positive solution we use (7) to obtain $\mathbf{H}_{2}$, which is then decomposed into two possible poses $(\mathbf{R}_{2},\mathbf{t}_{2})$. We then use these poses and focal lengths to triangulate three of the sampled points in the first two views thus obtaining points $\mathbf{X}_{i1}$. We use the corresponding points in the third view to obtain $\mathbf{R}_{3}$ and $\mathbf{t}_{3}$ using the $\mathbf{P}3\mathbf{P}$ solver [ding2023revisiting]. Note that it is possible to obtain $\mathbf{R}_{3}$ and $\mathbf{t}_{3}$ by decomposing $\mathbf{H}_{3}$ and using a single correspondence to obtain the scale of $\mathbf{t}_{3}$, but in practical experiments we found the approach using $\mathbf{P}3\mathbf{P}$ faster and more accurate. We score the generated models using pairwise Sampson error for each of the three image pairs. Whenever a new so-far-best model is found we perform local optimization using the Levenberg-Marquardt algorithm minimizing the pairwise Sampson error across the three pairs of views. In addition to the estimated focal lengths this strategy also produces the relative poses of all three cameras. We denote the proposed robust estimators as $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ for Case I, $\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ for Case II, $\mathbf{H}_{f\rho\rho}+\mathbf{P}3\mathbf{P}$ for Case III and $\mathbf{H}_{f\rho}+\mathbf{P}3\mathbf{P}$ for Case IV to distinguish them from the solvers. Our implementation is based on PoseLib [poselib].
@@ 6 Experiments
We perform extensive experiments on synthetic and real data to evaluate the performance of the proposed solvers and the robust focal length estimation strategy described in Section 5.3. We compare our solvers with several baselines using either pairwise or triplet correspondences. For pairwise correspondences, we consider the 6 point solver for relative pose with one unknown shared focal length [kukelova2012polynomial] denoted as $f\mathbf{E}f$and its combination with the $4+1$ point plane and parallax solver [torii2011six] using the DEGENSAC framework [chum2005degensac], which we denote as $f\mathbf{E}f+\mathbf{P}\mathbf{P}$. We also evaluate the solver for one unknown focal length [bujnak20093d] denoted as $\mathbf{E}f$. When considering triplets, we use a strategy [nister2004efficient] of using a pairwise solver followed by triangulation of points and registration of the third camera using the $\mathbf{P}3\mathbf{P}$ solver [ding2023revisiting]. (see Section 5.3 for more details). We denote the methods that work with triplet correspondences as $f\mathbf{E}f+\mathbf{P}3\mathbf{P}$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$, $\mathbf{E}f+\mathbf{P}3\mathbf{P}$ respectively.
Note that we do not compare with the solvers proposed in [heikkila2017using]. As visible from Table 1 these solvers are significantly slower than our solvers  [^2] The reported runtimes are runtimes of the original Matlab implementations of [heikkila2017using] and the Matlab + mex implementations of our solvers on I7-11700K CPU. and they return significantly more solutions that need to be tested inside RANSAC. As such they are not practical for real-world applications. Moreover, the solver for Case IV proposed in [heikkila2017using] returns solutions only to one of the two unknown focal lengths.
@@ 6.1 Synthetic experiments
[…]
@@
Case I
$\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ ours 4 triplets 0.0439 0.1491 51.49 65.58 49.31
$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$ [tzamos2023relative] 6 triplets 0.0503 0.2215 47.83 61.45 37.44
@@
$f\mathbf{E}f$ [kukelova2012polynomial] 6 pairs 0.4793 0.7143 10.49 17.07 24.54
$f\mathbf{E}f+\mathbf{P}\mathbf{P}$ [torii2011six] 6 pairs 0.5166 0.7935 10.31 16.51 31.95
Case II
$\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ ours 4 triplets 0.0611 0.2244 43.11 55.54 35.97
$f\mathbf{E}f+\mathbf{P}3\mathbf{P}$ [tzamos2023relative] 6 triplets 0.0692 0.2438 40.42 52.35 31.81
@@
$f\mathbf{E}f$ [kukelova2012polynomial] 6 pairs 0.4623 0.7085 12.01 19.12 28.59
$f\mathbf{E}f+\mathbf{P}\mathbf{P}$ [torii2011six] 6 pairs 0.4968 0.7875 12.11 18.98 37.60
$\mathbf{E}f$ [bujnak20093d] 6 pairs 0.6827 0.6843 18.87 13.01 22.59
@@
[…]
@@
The synthetic data is generated with $n=200$ 3D points visible by the three $1920\text{px}\times 1080\text{px}$ cameras with focal lengths sampled unfiormly $f_{g}\in\left[300,3000\right]$. The cameras are positioned randomly such that their views overlap. $n_{p}$ of the total $n$ points lie on a plane with random orientation. From the 3D points we create triplets with 75% inlier ratio by randomly shuffling 25% of the points. We add Gaussian noise with standard deviation of $\sigma$ to the pixel coordinates of points. We run all of the tested methods in PoseLib [poselib] for a fixed number of 100 iterations with the Sampson error threshold of 3 px. For methods that use only pairwise correspondences we consider only the first two views. For each configuration we generate 100 random scenes.
Fig. 3 shows the results of the synthetic experiments in which we vary both $\sigma$ and $n_{p}$. For Case II we also show how the error in the known focal length $\rho$ affects the accuracy of the methods (Fig. 3 (f)). When considering Case I, $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ shows significantly better accuracy than the remaining methods when a large portion of points lie on a single dominant plane (Fig. 3 (a)) even under high levels of noise (Fig. 3 (b)). This indicates a clear superiority of $\mathbf{H}_{fff}+\mathbf{P}3\mathbf{P}$ when dealing with images of planar scenes. Similarly, for Case II, $\mathbf{H}_{ff}+\mathbf{P}3\mathbf{P}$ also shows improved focal length estimation accuracy when considering planar scenes, although the improvement over competing methods is not as significant. For fully planar scenes, as expected, the two-view solvers $f\mathbf{E}f$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ and $\mathbf{E}f$ fail to generate correct solutions (Fig. 3 (b),(e),(f)). For scenes with few of the plane points, e.g., scenes with $n_{p}/n\in\left[0.95,0.9\right]$, the $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ solver, i.e., the solver where the $4+1$ plane+parallax solver is utilized within DEGENSAC with $f\mathbf{E}f$ [torii2011six], outperforms the $f\mathbf{E}f$ solver (Fig. 3 (a)). This is not surprising since the $f\mathbf{E}f+\mathbf{P}\mathbf{P}$ solver [torii2011six] was developed to handle scenes with dominant planes and some off-the plane points. The solvers that work with triplet correspondences, i.e. $f\mathbf{E}f+\mathbf{P}3\mathbf{P}$, $f\mathbf{E}f+\mathbf{P}\mathbf{P}+\mathbf{P}3\mathbf{P}$, $\mathbf{E}f+\mathbf{P}3\mathbf{P}$, outperform the two-view solvers.
@@ 6.2 Real-World Experiments
Since there are no suitable datasets for evaluating focal length recovery on planar scenes, we have collected a new dataset for evaluation which we will make publicly available. We have collected the dataset using 14 different cameras. For each camera we obtained ground truth focal lengths using the standard calibration method [zhang2000calibration] with a checkerboard pattern. Using these cameras we have captured four indoor and two outdoor planar scenes from various positions resulting 1870 images in total. Sample images of the scenes are shown in Fig. 5. To obtain triplet correspondences we used SuperPoint features [detone2018superpoint] with LightGlue matches [lindenberger2023lightglue]. More details are available in SM.
@@
Using the proposed solvers we may obtain multiple solutions. Therefore these solvers need to be used within a robust estimation framework such as RANSAC. Other simpler strategies such as voting [bujnak2009robust] may fail due to the fact that more than one solution may be geometrically feasible.
Moreover, as mentioned above for Case III and IV focal lengths can not be distinguished using only planar points thus requiring some off-plane points during scoring. Note that this is a property of the problem and not the solvers. Problem with recovering one focal length for Case III was mentioned also in [heikkila2017using].
@@ 8 New Constraints
Here we provide details on the constraints introduced in Sec. 4 of the main paper. We show the steps for their derivation including Macaulay2 [M2] code.
To derive the constraints relating the focal lengths and the elements of the matrix $\mathbf{Q}_{i}$, we first create an ideal $I$ [cox2005using] generated by the 12 polynomials extracted form the matrix equation Eq (16) in the main paper, i.e. the equation
[…]
@@
In the next step, the unknown elements of the normal vector $\mathbf{n}$, i.e., $n_{x},n_{y},n_{z}$, and the scale factors $s_{2},s_{3}$ are eliminated from the generators of $I$ by computing the generators of the elimination ideal $J=I\cap\mathbb{C}[q_{21},\dots,q_{36}]$. Here, $q_{j.}$ are the entries of $\mathbf{Q}_{2},\mathbf{Q}_{3}$. These generators can be computed, for example, in the computer algebra software Macaulay2 [M2] using the following code:
@@
[…]
@@ 10 Evaluation for Case III and Case IV
In this section we evaluate the robust estimators $\mathbf{H}_{\rho ff}+\mathbf{P}3\mathbf{P}$ for Case III and $\mathbf{H}_{f\rho}+\mathbf{P}3\mathbf{P}$ for Case IV which were presented in Sec.5.3 of the main paper. As baselines for comparison we use the combination of the $f\mathbf{E}f$solver [kukelova2012polynomial] or its DEGANSAC [chum2005degensac] variant with the plane and parallax solver [torii2011six] together with the $\mathbf{P}4\mathbf{P}f$ solver [kukelova2016efficient] for Case III and the $\mathbf{E}f$ solver in combination with $\mathbf{P}4\mathbf{P}f$ solver [kukelova2016efficient] for Case IV.
@@ 10.1 Multiple Geometrically Valid Solutions
We observe that for both Case III and Case IV a single planar scene may result in multiple geometrically valid solutions. This makes it difficult to select a single correct solution for a given set of point correspondences. This leads to generally worse performance of the proposed solvers than for Case I and Case II. Note, that this is a feature of the problems and not the solvers. Problem with recovering one focal length for Case III was mentioned also in [heikkila2017using].
@@ 10.3 Real-World Experiments
In Tab. 3 we show the results comparing our method and the baseline approaches in three different variants. As the first variant we do not discard any solutions. For the second variant, we set range of acceptable FOVs by considering the prior for focal lengths used by the popular SfM software COLMAP [schonberger2016structure] which is set as $f_{p}=1.2\text{max}(width,height)$, which corresponds to a field of view of $\sim 45^{\circ}$. To obtain a range we consider 30% increase or decrease in focal length, resulting in the acceptable field of view range from 35.5 to 61.5. As the last variant we use a range of 50 to 70. We chose this range since most commercially available phone cameras fall within it.
All variants were implemented in PoseLib [poselib]. We set the maximum epipolar threshold to 3 px. We used early termination with 0.9999 confidence and a minimum of 100 iterations and a maximum of 1000.